satellite73 (satellite73):

Factor \[a^4+2a^2b^2+9b^4\]

OpenStudy (anonymous):

a^4 + 2a^2b^2 + 9b^4 = a^4 + 6a^2b^2 + 9b^4 - 4a^2b^2 = (a^2 + 3b^2)^2 - (2ab)^2 = (a^2 + 2ab + 3b^2) (a^2 - 2ab + 3b^2).

OpenStudy (nincompoop):

\[(a^2-2 a b+3 b^2) (a^2+2 a b+3 b^2)\]

OpenStudy (anonymous):

guess i have to come up with a harder one

OpenStudy (anonymous):

ahah :p

OpenStudy (nincompoop):

try something with >2 leading coefficient

OpenStudy (xapproachesinfinity):

you want harder ones?